OpenStudy (anonymous):

evaluate {cos x} tan^-1(1/2) = x = 0

OpenStudy (anonymous):

\[\cos (x)\tan^{-1} (\frac{ 1 }{ 2 })=x=0\]

OpenStudy (anonymous):

is that your question?

OpenStudy (anonymous):

cos(x) [tan−1(1/2)]=x=0 what is cos(tan−1(1/2)) I know cos(0) =1